What is the value or values of n in the equation 2n2-7n+6=0
Answer by ewatrrr(24785)


Hi,
solving for n
2n^2-7n+6=0
factoring
(2n - 3)(n -2)= 0 Note:SUM of the inner product(-3n) and the outer product(-4n) = -7n
(2n - 3)=0
n = 3/2
(n -2)= 0
n = 2